# Eaman Jahani

### v6

@@ 1 Introduction
Generative AI is being integrated into work practices across the economy (zhang2023generative, bright2024generative), yielding notable productivity gains in tasks as diverse as software development, writing, and scientific research (brynjolfsson2023generative, dell2023navigating, Noy2023, peng2023impact, Yu2024AIImpact). Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning2024Automated), including tasks as complex as chemical research and proving mathematical theorems (RomeraParedes2024Math, Boiko2023AutonomousResearch). The adoption of generative AI is also occurring at an unprecedented pace, with recent research showing that approximately 28% of U.S. workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (bright2024generative, bick2024rapid).
@@
As with many other general-purpose technologies, the effectiveness of generative AI depends not only on the technology itself, but on users’ ability to craft inputs that produce high-quality results. To interact with generative AI systems, users provide written instructions—or prompts—that guide the model’s behavior. These prompts can range from simple commands (e.g., “write a short story about a robot”) to highly detailed specifications tailored to particular outputs (e.g., a series of paragraphs instructing an AI system to implement a complete piece of software). In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (brynjolfsson2000beyond).
@@
Prompting has quickly become an area of active research and practice. Scholars have developed taxonomies of prompt engineering techniques (oppenlaender2023), documented recurring patterns in prompt construction (schulhoff2024prompt), and examined how developers embed prompts into software systems (liang2024prompts). Other studies have explored prompting strategies for specific applications, including image generation (donyehiya2023, xie2023) and clinical documentation (yao2024clinicians). In parallel, practitioners have built prompt libraries, shared tutorials, and developed tools to support prompt design. These developments signal a growing consensus that prompting plays a meaningful role in extracting value from generative AI systems.
@@
Yet despite this consensus, prompting remains understudied as a dynamic practice. Many prompt libraries and tutorials present effective prompts as reusable artifacts. But prompts that work well with one model version may underperform or break entirely with the next (liang2024prompts, meincke2025prompting). While recent research increasingly views prompting as an adaptive process, empirical evidence remains limited on how these strategies evolve—both as users refine prompts for a single model and as they adjust to model updates—and on how these changes ultimately affect performance. This raises a broader question for individuals and organizations investing in prompting capabilities: Are prompt strategies transferable across model versions, or must they be continually revised to match changing model behavior?
@@
Generative AI systems support a wide variety of task structures. Some tasks are steering-oriented, where users aim to reach a clearly defined goal and success, can be measured against an objective benchmark (e.g., write code to solve a specific problem). Others are creative, where goals are open-ended, evaluation is subjective, and improvement depends more on exploration and curation than on precision (e.g., make a beautiful picture). Recent work helps to establish these regimes: vafa2025s investigate the steerability of models, i.e., how effectively users can reach a specified target, whereas orwig2024language and zhou2024generative describe generative AI–assisted creation as a co-creative process of idea generation, filtering, and aesthetic evaluation. In organizational settings, steering tasks (or bounded tasks, herein) dominate operational workflows such as classification, brand-compliant content generation, or reproducing structured materials where quality can be defined ex ante. Creative tasks (or unbounded tasks, herein), by contrast, underpin innovation activities such as logo or campaign design, product ideation, and concept development, where success depends on novelty, appeal, and fit to evolving goals. Understanding how prompting and prompt adaptation operate across these two regimes is essential for organizations deploying generative AI at scale.
@@
In terms of related literature and additional theory, our research builds on work in information systems, emphasizing the importance of dynamic, user-driven complements to digital technologies. Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (bharadwaj2000resource, joshi2010changing, teece1997dynamic). Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (jasperson2005comprehensive). Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (fugener2022cognitive). And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (bohm2023co-evolution).
We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (brynjolfsson1993productivity, brynjolfsson2000beyond, brynjolfsson2021productivity, david1990dynamo). We conceptualize prompt adaptation as a dynamic complement that co-evolves with model capability, emerges through use rather than formal training, and whose importance depends on task structure and system design. In this way, prompt adaptation shapes how—and under what conditions—technical improvements translate into downstream economic value.
@@ 2.1 Notation and Problem Setting
Let $\theta\in(0,1]$ denote the model’s capacity to translate prompts into high-fidelity outputs, $s\in(0,1]$ denote a user’s baseline skill in prompt engineering, and $x\geq 0$ denote the effort the user expends on writing and refining prompts. Each unit of effort incurs a linear cost $c(x)=kx$ for some $k>0$, reflecting the total cognitive and temporal costs of refining one’s prompts. We assume $k$ is small ($k\!\approx\!0$), consistent with the low marginal cost of iterative prompting and using generative AI models more generally (shahidi2025coasean).
@@ 3.2.1 Outcome Definition: Expected CLIP Embedding Cosine Similarity
The primary outcome in the image replication experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (radford2021learning). CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together. By embedding both the target image and each generated image into this space and computing the cosine similarity between them, we obtain a quantitative measure of how closely the generated image matches the target along both visual and conceptual dimensions.
Because the output of each generative model is stochastic, the same prompt can yield different images across attempts. To account for this variability, we generated 10 images for each prompt and computed their cosine similarity to the target image individually. We then averaged these 10 similarity scores to produce an expected quality score for each prompt—the primary outcome variable used in our analysis. As a robustness check, we replicated all analyses using DreamSim, a recently developed alternative to cosine similarity on CLIP embeddings that is based on perceptual similarity and aligns more closely with human judgments. [^8] DreamSim (fu2023dreamsim) is designed to better capture human perceptions of image similarity than traditional embedding-based methods. Our results are robust to using DreamSim in place of CLIP cosine similarity. Full DreamSim-based analyses can be found in Appendix H. The two measures were highly correlated, and our findings were consistent across both.
@@ 3.3.1 Outcome Definition: Bradley-Terry Strength
Because each logo-generation task was open-ended and lacked an objective reference image, we could not assess output quality using embedding-based similarity metrics (e.g., cosine similarity of CLIP representations). Instead, we employed the Bradley–Terry (BT) model (bradley1952rank), a standard approach for inferring latent quality scores when only relative judgments are available (zermelo1929berechnung, ford1957solution). At a high level, the BT model is a probabilistic model that estimates a strength parameter (which we refer to hereafter as “BT $\beta$”) for each item based on the outcomes of many pairwise comparisons.
@@ 3.4 Replay Analysis for Separating Model and Prompting Effects
In the image replication experiment, replay was conducted several weeks after the initial data collection, so we additionally re-generated outputs for the original prompt–model combinations contemporaneously with replay to guard against model drift (chen2023gptchanging). In the logo generation experiment, replay occurred shortly after data collection, so this additional step was unnecessary.
@@ 5 Discussion
A related question we address is whether prompt adaptation must occur through direct user experimentation or whether it can be partially automated by the system itself. In practice, many generative AI systems implement automated prompt rewriting to reduce user effort, implicitly positioning automation as a substitute for user-driven adaptation. Our findings suggest that the effectiveness of such automation depends critically on task structure and alignment with user goals. In the bounded image replication task, the DALL-E 3 endpoint’s automated prompt rewriting reduced realized gains, consistent with misalignment between the rewriting logic and users’ objective of faithfully reproducing a target image. By contrast, in the open-ended logo generation task, the same rewriting mechanism modestly improved performance. Together, these findings underscore the importance of careful design in automated prompt rewriting systems. When system prompts are well aligned with user goals and task structure, automated rewriting can improve performance. but poorly aligned system prompts, especially when hidden, can degrade performance and/or limit users’ ability to discover or exploit emergent but useful ways of interacting with a generative AI system. Such complexities are highlighted by related work, such as yao2024clinicians, which studies automated prompt optimization in the context of clinical documentation.
@@
More generally, our results both align with and help clarify themes in the literature on human–AI collaboration and the economics of technology, particularly the role of dynamic complements such as prompt adaptation that co-evolve with technical change (bohm2023co-evolution). As generative AI models advance, often substantially from month to month, organizations that fail to adapt their prompting strategies in task settings where precision and steerability matter may forgo a meaningful share of the economic value these upgrades make possible. This co-evolutionary pattern is not unique to AI. It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (brynjolfsson1993productivity, david1990dynamo). However, the pace of generative AI advancement introduces a distinctive challenge: the window for adaptation is far shorter. Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (brynjolfsson2000beyond, brynjolfsson2021productivity). Our findings complement this stream of work by offering direct empirical evidence that a specific dynamic complement—user adaptation, specifically through prompt refinement—can account for a substantial share of realized performance gains from model upgrades in task settings where precision and steerability matter.
@@
Seen from a broader organizational perspective, these findings also speak to long-standing work on IT-enabled dynamic capabilities (bharadwaj2000resource, joshi2010changing, teece1997dynamic) and post-adoptive IT use (jasperson2005comprehensive). Prompt adaptation, as we observe it, is not grounded in formal training or specialized skill. Participants in our study, drawn from a general population rather than a specialized group of prompt engineers, improved performance through trial-and-error within a single session. This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (attewell1992technology, von2006democratizing). The relative accessibility of prompt adaptation highlights one potential pathway toward more broadly distributed productivity gains, provided that users are supported by scaffolds and interfaces that enable iterative refinement (rogers2003diffusion). At the same time, this accessibility introduces risk. Over-optimizing for a particular model version may reduce users’ ability to adapt as systems evolve. This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (henderson1990architectural). Supporting long-term adaptation may require not just prompt training, but workflows and learning mechanisms that encourage ongoing experimentation.
@@
For researchers and practitioners interested in the economics of AI, our findings reinforce the idea that complements play a central role in shaping performance. In some task settings, prompt adaptation can account for as much as half of the realized performance gains from model upgrades, underscoring that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (arthur2009nature, dosi1982technological). As generative AI continues to advance, organizations that invest in adaptive, not static, complementary skills will be better positioned to realize its full value.
@@ D.1 Image Similarity
We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (fu2023dreamsim). In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers. Our results are qualitatively and quantitatively similar using DreamSim instead.
@@
To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (radford2021learning) using the version of CLIP available on Hugging Face (CLIP-embeddings) for each participant-generated image and for each target image. Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text. We then calculated the cosine similarity between each participant-generated image’s CLIP embedding and the relevant target image’s CLIP embedding.
@@
DreamSim is an image similarity measure proposed recently by (fu2023dreamsim). The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity. Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score. After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.

### v7

@@ 1 Introduction
Generative AI is being integrated into work practices across the economy (Zhang and Kamel Boulos 2023, Bright et al. 2024), yielding notable productivity gains in tasks as diverse as software development, writing, and scientific research (Brynjolfsson et al. 2023, Dell’Acqua et al. 2023, Noy and Zhang 2023, Peng et al. 2023, Yu 2024). Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning et al. 2024), including tasks as complex as chemical research and proving mathematical theorems (Romera-Paredes et al. 2024, Boiko et al. 2023). The adoption of generative AI is also occurring at an unprecedented pace, with recent research showing that approximately 28% of U.S. workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (Bright et al. 2024, Bick et al. 2024).
@@
As with many other general-purpose technologies, the effectiveness of generative AI depends not only on the technology itself, but on users’ ability to craft inputs that produce high-quality results. To interact with generative AI systems, users provide written instructions—or prompts—that guide the model’s behavior. These prompts can range from simple commands (e.g., “write a short story about a robot”) to highly detailed specifications tailored to particular outputs (e.g., a series of paragraphs instructing an AI system to implement a complete piece of software). In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (Brynjolfsson and Hitt 2000).
@@
Prompting has quickly become an area of active research and practice. Scholars have developed taxonomies of prompt engineering techniques (Oppenlaender 2023), documented recurring patterns in prompt construction (Schulhoff et al. 2024), and examined how developers embed prompts into software systems (Liang et al. 2024). Other studies have explored prompting strategies for specific applications, including image generation (Don-Yehiya et al. 2023, Xie et al. 2023) and clinical documentation (Yao et al. 2024). In parallel, practitioners have built prompt libraries, shared tutorials, and developed tools to support prompt design. These developments signal a growing consensus that prompting plays a meaningful role in extracting value from generative AI systems.
@@
Yet despite this consensus, prompting remains understudied as a dynamic practice. Many prompt libraries and tutorials present effective prompts as reusable artifacts. But prompts that work well with one model version may underperform or break entirely with the next (Liang et al. 2024, Meincke et al. 2025). While recent research increasingly views prompting as an adaptive process, empirical evidence remains limited on how these strategies evolve—both as users refine prompts for a single model and as they adjust to model updates—and on how these changes ultimately affect performance. This raises a broader question for individuals and organizations investing in prompting capabilities: Are prompt strategies transferable across model versions, or must they be continually revised to match changing model behavior?
@@
Generative AI systems support a wide variety of task structures. Some tasks are steering-oriented, where users aim to reach a clearly defined goal and success, can be measured against an objective benchmark (e.g., write code to solve a specific problem). Others are creative, where goals are open-ended, evaluation is subjective, and improvement depends more on exploration and curation than on precision (e.g., make a beautiful picture). Recent work helps to establish these regimes: Vafa et al. (2025) investigate the steerability of models, i.e., how effectively users can reach a specified target, whereas Orwig et al. (2024) and Zhou and Lee (2024) describe generative AI–assisted creation as a co-creative process of idea generation, filtering, and aesthetic evaluation. In organizational settings, steering tasks (or bounded tasks, herein) dominate operational workflows such as classification, brand-compliant content generation, or reproducing structured materials where quality can be defined ex ante. Creative tasks (or unbounded tasks, herein), by contrast, underpin innovation activities such as logo or campaign design, product ideation, and concept development, where success depends on novelty, appeal, and fit to evolving goals. Understanding how prompting and prompt adaptation operate across these two regimes is essential for organizations deploying generative AI at scale.
@@
In terms of related literature and additional theory, our research builds on work in information systems, emphasizing the importance of dynamic, user-driven complements to digital technologies. Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (Bharadwaj 2000, Joshi et al. 2010, Teece et al. 1997). Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (Jasperson et al. 2005). Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (Fügener et al. 2022). And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (Böhm and Schedlberger 2023).
We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (Brynjolfsson 1993, Brynjolfsson and Hitt 2000, Brynjolfsson et al. 2021, David 1990). We conceptualize prompt adaptation as a dynamic complement that co-evolves with model capability, emerges through use rather than formal training, and whose importance depends on task structure and system design. In this way, prompt adaptation shapes how—and under what conditions—technical improvements translate into downstream economic value.
@@ 2.1 Notation and Problem Setting
Let $\theta\in(0,1]$ denote the model’s capacity to translate prompts into high-fidelity outputs, $s\in(0,1]$ denote a user’s baseline skill in prompt engineering, and $x\geq 0$ denote the effort the user expends on writing and refining prompts. Each unit of effort incurs a linear cost $c(x)=kx$ for some $k>0$, reflecting the total cognitive and temporal costs of refining one’s prompts. We assume $k$ is small ($k\!\approx\!0$), consistent with the low marginal cost of iterative prompting and using generative AI models more generally (Shahidi et al. 2025).
@@ 3.2.1 Outcome Definition: Expected CLIP Embedding Cosine Similarity
The primary outcome in the image replication experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (Radford et al. 2021). CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together. By embedding both the target image and each generated image into this space and computing the cosine similarity between them, we obtain a quantitative measure of how closely the generated image matches the target along both visual and conceptual dimensions.
Because the output of each generative model is stochastic, the same prompt can yield different images across attempts. To account for this variability, we generated 10 images for each prompt and computed their cosine similarity to the target image individually. We then averaged these 10 similarity scores to produce an expected quality score for each prompt—the primary outcome variable used in our analysis. As a robustness check, we replicated all analyses using DreamSim, a recently developed alternative to cosine similarity on CLIP embeddings that is based on perceptual similarity and aligns more closely with human judgments. [^8] DreamSim (Fu et al. 2023) is designed to better capture human perceptions of image similarity than traditional embedding-based methods. Our results are robust to using DreamSim in place of CLIP cosine similarity. Full DreamSim-based analyses can be found in Appendix H. The two measures were highly correlated, and our findings were consistent across both.
@@ 3.3.1 Outcome Definition: Bradley-Terry Strength
Because each logo-generation task was open-ended and lacked an objective reference image, we could not assess output quality using embedding-based similarity metrics (e.g., cosine similarity of CLIP representations). Instead, we employed the Bradley–Terry (BT) model (Bradley and Terry 1952), a standard approach for inferring latent quality scores when only relative judgments are available (Zermelo 1929, Ford 1957). At a high level, the BT model is a probabilistic model that estimates a strength parameter (which we refer to hereafter as “BT $\beta$”) for each item based on the outcomes of many pairwise comparisons.
@@ 3.4 Replay Analysis for Separating Model and Prompting Effects
In the image replication experiment, replay was conducted several weeks after the initial data collection, so we additionally re-generated outputs for the original prompt–model combinations contemporaneously with replay to guard against model drift (Chen et al. 2023). In the logo generation experiment, replay occurred shortly after data collection, so this additional step was unnecessary.
@@ 5 Discussion
A related question we address is whether prompt adaptation must occur through direct user experimentation or whether it can be partially automated by the system itself. In practice, many generative AI systems implement automated prompt rewriting to reduce user effort, implicitly positioning automation as a substitute for user-driven adaptation. Our findings suggest that the effectiveness of such automation depends critically on task structure and alignment with user goals. In the bounded image replication task, the DALL-E 3 endpoint’s automated prompt rewriting reduced realized gains, consistent with misalignment between the rewriting logic and users’ objective of faithfully reproducing a target image. By contrast, in the open-ended logo generation task, the same rewriting mechanism modestly improved performance. Together, these findings underscore the importance of careful design in automated prompt rewriting systems. When system prompts are well aligned with user goals and task structure, automated rewriting can improve performance. but poorly aligned system prompts, especially when hidden, can degrade performance and/or limit users’ ability to discover or exploit emergent but useful ways of interacting with a generative AI system. Such complexities are highlighted by related work, such as Yao et al. (2024), which studies automated prompt optimization in the context of clinical documentation.
@@
More generally, our results both align with and help clarify themes in the literature on human–AI collaboration and the economics of technology, particularly the role of dynamic complements such as prompt adaptation that co-evolve with technical change (Böhm and Schedlberger 2023). As generative AI models advance, often substantially from month to month, organizations that fail to adapt their prompting strategies in task settings where precision and steerability matter may forgo a meaningful share of the economic value these upgrades make possible. This co-evolutionary pattern is not unique to AI. It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (Brynjolfsson 1993, David 1990). However, the pace of generative AI advancement introduces a distinctive challenge: the window for adaptation is far shorter. Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (Brynjolfsson and Hitt 2000, Brynjolfsson et al. 2021). Our findings complement this stream of work by offering direct empirical evidence that a specific dynamic complement—user adaptation, specifically through prompt refinement—can account for a substantial share of realized performance gains from model upgrades in task settings where precision and steerability matter.
@@
Seen from a broader organizational perspective, these findings also speak to long-standing work on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al. 2010, Teece et al. 1997) and post-adoptive IT use (Jasperson et al. 2005). Prompt adaptation, as we observe it, is not grounded in formal training or specialized skill. Participants in our study, drawn from a general population rather than a specialized group of prompt engineers, improved performance through trial-and-error within a single session. This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (Attewell 1992, Von Hippel 2006). The relative accessibility of prompt adaptation highlights one potential pathway toward more broadly distributed productivity gains, provided that users are supported by scaffolds and interfaces that enable iterative refinement (Rogers 2003). At the same time, this accessibility introduces risk. Over-optimizing for a particular model version may reduce users’ ability to adapt as systems evolve. This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (Henderson and Clark 1990). Supporting long-term adaptation may require not just prompt training, but workflows and learning mechanisms that encourage ongoing experimentation.
@@
For researchers and practitioners interested in the economics of AI, our findings reinforce the idea that complements play a central role in shaping performance. In some task settings, prompt adaptation can account for as much as half of the realized performance gains from model upgrades, underscoring that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (Arthur 2009, Dosi 1982). As generative AI continues to advance, organizations that invest in adaptive, not static, complementary skills will be better positioned to realize its full value.
@@ D.1 Image Similarity
We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (Fu et al. 2023). In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers. Our results are qualitatively and quantitatively similar using DreamSim instead.
@@
To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (Radford et al. 2021) using the version of CLIP available on Hugging Face (OpenAI 2024) for each participant-generated image and for each target image. Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text. We then calculated the cosine similarity between each participant-generated image’s CLIP embedding and the relevant target image’s CLIP embedding.
@@
DreamSim is an image similarity measure proposed recently by (Fu et al. 2023). The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity. Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score. After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.